Here's a scenario our professor asked us to figure out for extra credit and I am not sure how to set it up or solve:

Jeremy has a 64.80% in his class and needs a 70% in order to pass. The grading rubric is made up of three tests equaling 20% of his grade, online homework, which is 15% of his grade, and final exam, which is 25% of the overall grade.

Jeremy's three test scores are 68%, 40%, and 60% and has gotten a 100% on all of the homework, which has him currently sitting at a 64.80%.

The professor told her students that she will take their lowest test grade and replace it with their final grade (if it is a higher score) while also keeping the final grade (meaning two of their test grades will be the same).

For example, if the student got a 70% on the final, his test grades would change to 68%, 70% (replacing the 40%), 60% (weighted 20%) and then 70% for final (weighted 25%).

What does Jeremy need to get on the final exam in order for his second exam (40% weighted at 20% of his grade) to be replaced, while also including that same grade as his final (weighted at 25%) in order to get a 70% overall in the class?
Answer #1

ANSWER:

Let us represent the scores as followes

= score in test 1

= score in test 2

= score in test 3

h = score in test home work

f = score in test final exam.

Given weightage for = = 20%= 20/100

h = 15/100 = 15%

f = 25/100 = 25%

Now , A/Q

Total final score   70%

Total final score 70 / 100

Now , total final score on the given weightage is calculated as

But it is given in the question   = f

  

  

Minimum final grade should be 66% in order to get an overall grade of 70%.